# Exploring Test-Time Adaptation for Object Detection in Continually Changing Environments

### v1

@@ Abstract.
For real-world applications, neural network models are commonly deployed in dynamic environments, where the distribution of the target domain undergoes temporal changes. Continual Test-Time Adaptation (CTTA) has recently emerged as a promising technique to gradually adapt a source-trained model to test data drawn from a continually changing target domain. Despite recent advancements in addressing CTTA, two critical issues remain: 1) The use of a fixed threshold for pseudo-labeling in existing methodologies leads to the generation of low-quality pseudo-labels, as model confidence varies across categories and domains; 2) While current solutions utilize stochastic parameter restoration to mitigate catastrophic forgetting, their capacity to preserve critical information is undermined by its intrinsic randomness. To tackle these challenges, we present CTAOD, aiming to enhance the performance of detection models in CTTA scenarios. Inspired by prior CTTA works for effective adaptation, CTAOD is founded on the mean-teacher framework, characterized by three core components. Firstly, the object-level contrastive learning module tailored for object detection extracts object-level features using the teacher’s region of interest features and optimizes them through contrastive learning. Secondly, the dynamic threshold strategy updates the category-specific threshold based on predicted confidence scores to improve the quality of pseudo-labels. Lastly, we design a data-driven stochastic restoration mechanism to selectively reset inactive parameters using the gradients as weights for a random mask matrix, thereby ensuring the retention of essential knowledge. We demonstrate the effectiveness of our approach on four CTTA tasks for object detection, where CTAOD outperforms existing methods, especially achieving a 3.0 mAP improvement on the Cityscapes-to-Cityscapes-C CTTA task. The code of this work will be released soon.
@@ 1. Introduction
Figure 1. Motivation of our proposed method. The source-trained model adapts to dynamic environments without source data. (1) DT: The bar chart above represents the mean model scores for different categories across four domains of the ACDC dataset (Sakaridis et al., 2021). While existing methods produce the pseudo-labels via a fixed threshold, CTAOD utilizes DT to update the category-specific threshold based on the mean predicted scores. (2) DSR: The bottom describes the neuron reset process, where neurons with deeper shades of purple indicate increased activity levels and blue neurons represent corresponding neurons from the source model. While existing methods restore parameters randomly, CTAOD resets inactive parameters with higher possibilities.
@@
Deep learning models have demonstrated immense potential across various data modalities such as image (He et al., 2016; Dosovitskiy et al., 2020; Ren et al., 2015), audio (Aytar et al., 2016; Amodei et al., 2016), language (Devlin et al., 2018; Vaswani et al., 2017), and video (Tran et al., 2015; Carreira and Zisserman, 2017). However, these models experience pronounced degradation in performance when confronted with training data (i.e., source domain) and testing data (i.e., target domain) originating from disparate distributions. This phenomenon, commonly referred to as distribution shifts, poses a significant challenge (Geirhos et al., 2018; Hendrycks and Dietterich, 2019; Luo et al., 2020; Deng et al., 2014). In such a scenario, unsupervised domain adaptation (UDA) becomes crucial, which typically involves aligning the source and target data distributions, thereby mitigating the impact of distribution shifts (Ganin et al., 2016; Tzeng et al., 2017; Saito et al., 2018). Still, UDA falls short by necessitating access to source data, which is often inaccessible due to privacy constraints or data transmission barriers (VS et al., 2023b; Huang et al., 2021).
@@
This limitation catalyzes the exploration of Test-Time Adaptation (TTA) where the source-trained model directly adapts toward unlabeled test samples encountered during evaluation in an online manner, without the reliance on the source data (Mummadi et al., 2021; Sinha et al., 2023; You et al., 2021; Wang et al., 2020). Nonetheless, the aforementioned methods, which assume a static target domain, face a more challenging and realistic problem, as real-world machine learning systems work in non-stationary and continually evolving environments (Wang et al., 2022; Döbler et al., 2023). Existing TTA methods are vulnerable to catastrophic forgetting of previously learned source knowledge and error accumulation when adaptation faces more than one distribution shift (Wang et al., 2022; Döbler et al., 2023; Niloy et al., 2024). For example, a vehicle may encounter various continuous environmental changes such as fog, night, rain, and snow during its journey.
@@
Recently, Wang et al. (2022) introduce CoTTA by applying stochastic parameters restoration to mitigate catastrophic forgetting in Continual Test-Time Adaptation (CTTA) scenarios (see Table 1), where the model is continually adapted to sequences of target domains. Although the strategy of randomly selecting parameters for restoration partially helps alleviate the forgetting of source knowledge (Wang et al., 2022; Zhu et al., 2023), its randomness may also contribute to the loss of crucial knowledge specific to the current domain (refer to Figure 1). Additionally, existing CTTA methods normally rely on self-training methods by utilizing pseudo-labels through a fixed threshold to supervise the training of the model (Wang et al., 2022; Döbler et al., 2023; Sójka et al., 2023). Given that model confidence could vary across categories and domains (refer to Figure 1), employing a uniformly fixed threshold could result in excluding high-quality pseudo-labels while incorporating incorrect ones. Erroneous or missing pseudo-labels lead to error accumulation as they provide negative feedback to the model.
@@
Therefore, despite recent efforts to handle CTTA, two significant challenges persist: Current self-training-based methods suffer from noisy and low-quality pseudo-labels, leading to error accumulation (Challenge 1); Continual adaptation to dynamic environments in existing methods struggles to effectively retain valuable knowledge about the current domain while mitigating noise specific to the previous domain (Challenge 2).
@@
To tackle the identified challenges in the CTTA setting for object detection, we introduce CTAOD (Continual Test-time Adaption for Object Detection). Aligning with previous CTTA works for robust adaptation (Wang et al., 2022; Gong et al., 2022; Döbler et al., 2023), CTAOD is constructed upon the mean-teacher framework (Tarvainen and Valpola, 2017) where the teacher model is an exponential moving average of the student model. Particularly, CTAOD comprises Object-level Contrastive Learning (OCL), Dynamic Threshold (DT), and Data-driven Stochastic Restoration (DSR).
@@
More specifically, we propose OCL, which is tailored for object detection, to acquire fine-grained and localized feature representation, motivated by the effectiveness of contrastive learning (CL) in self-supervised learning (Chen et al., 2020; He et al., 2020; Grill et al., 2020; Chen and He, 2021). OCL extracts features based on proposals generated by the Region Proposal Network (RPN), which provides multiple cropped views around the object instance at different locations and scales. Subsequently, CL loss is applied on RPN cropped views to guide the model to encourage similar object instances to remain close while pushing dissimilar ones apart. The OCL is well integrated into the mean-teacher paradigm as a drop-in enhancement for feature adaptation. Furthermore, to improve the quality of pseudo-labels across various categories, i.e., addressing Challenge 1, we devise a DT strategy to dynamically adjust the threshold for each category individually based on the predicted confidence scores. The dynamic nature of the DT method makes it better suited to address the effects of continuously changing distributions compared to fixed thresholds. Finally, the DSR mechanism is utilized to reset inactive parameters with a higher possibility than active ones by employing the gradients of parameters as weights for a random mask matrix. On the one hand, it not only helps alleviate the impact of noise from the previous domain but also preserves important information, i.e., addressing Challenge 2. On the other hand, its randomness also allows potentially falsely activated parameters to be reset, thereby making the model more stable.
@@
We demonstrate the effectiveness of our proposed approach on four continual test-time adaptation benchmark tasks for object detection. These datasets are associated with synthetic and real-world distribution shifts in the short-term and long-term adaption (i.e. Cityscapes (Cordts et al., 2016) $\rightarrow$ Cityscapes-C (Hendrycks and Dietterich, 2019), SHIFT (Sun et al., 2022), and Cityscapes (Cordts et al., 2016) $\rightarrow$ ACDC (Sakaridis et al., 2021) adaptation). Experimental results indicate that our method significantly improves performance over existing state-of-the-art methods, with gains of up to 3.0 mAP on these tasks.
In summary, the main contributions of this paper are as follows:
@@
  • This study introduces a novel method named CTAOD, which pioneers exploring CTTA for detection models. Specifically, we propose to leverage object-level features for contrastive learning to refine feature representation for object detection.
@@
  • To address two challenges in CTTA, we dynamically update the category-specific threshold based on predicted scores to improve the quality of pseudo-labels and reset the inactive parameter with higher possibilities to mitigate forgetting.
@@
  • Empirical experiments demonstrate that our proposed method surpasses existing methods and effectively facilitates short-term and long-term adaptation in dynamic environments.
@@ 2.1. Source-Free Domain Adaptation
UDA tackles the inter-domain divergence by aligning the distributions of source and target data (Ganin et al., 2016; Li et al., 2019; Luo et al., 2020; Jiang et al., 2020). Despite its effectiveness, the limitation of UDA lies in its requirement for access to the source domain data, which often raises concerns regarding data privacy, data portability, and data transmission efficiency. As a result, Source-Free Domain Adaptation (SFDA) received extensive research attention, especially in object detection, where the source-trained detector is re-trained on the target training data (VS et al., 2023b, a; Li et al., 2022b, 2021a, 2021b) before evaluation. For instance, MemCLR (VS et al., 2023a) employs a cross-attention transformer-based memory bank for contrastive learning, while IRG (VS et al., 2023a) utilizes the object relations with instance relation graph network to explore the SFDA setting for object detection without the source data. However, the standard SFDA setting requires prior knowledge of the target domain, which is impractical in most real-world applications.
@@ 2.2. Test-Time Adaptation
TTA adapts the source-trained model to the target test data during inference time without access to the source data. Since both TTA and SFDA involve adapting the source-trained model to the unlabeled target data without utilizing source data, some works also refer to TTA as SFDA (Wang et al., 2022; Brahma and Rai, 2023). In this paper, we distinguish between TTA and SFDA based on evaluation protocol, although they can be transformed into each other in experiments. Furthermore, TTA methods improve the model performance under distribution shift commonly through techniques like pseudo-labeling (Sinha et al., 2023; Iwasawa and Matsuo, 2021; Sun et al., 2020; Zeng et al., 2023), batchnorm statistics updating (Hu et al., 2021; You et al., 2021), or entropy regularization (Wang et al., 2020; Iwasawa and Matsuo, 2021; Niu et al., 2022) during testing. For example, Tent (Wang et al., 2020) updates the batchnorm parameters with entropy minimization and demands a large batch size for optimization during test-time adaptation, which is unsuitable for real-time detection model deployment where images are processed sequentially. The above approaches assume a specific and static target domain where all the test inputs come from a single domain. However, in numerous practical scenarios, the distribution of test input may exhibit a continual shift over time.
@@ 2.3. Continual Test-time Adaptation
Early works consider adaptation to evolving and continually changing domains by aligning the source and target data (Hoffman et al., 2014; Wulfmeier et al., 2018). These methods rely on source data during inference, which limits their applicability. A significant advancement in this area is the proposal of CoTTA (Wang et al., 2022), marking the first work tailored to the demands of CTTA by adapting a pre-trained model to sequences of domains without using any source data. Subsequently, research efforts have been dedicated to exploring CTTA, primarily focused on classification (Döbler et al., 2023; Niloy et al., 2024; Brahma and Rai, 2023; Yu et al., 2023) and segmentation (Niloy et al., 2024; Song et al., 2023; Zhu et al., 2023) tasks. For instance, similar to CoTTA (Wang et al., 2022), Zhu et al. (2023) apply a stochastic reset mechanism for CTTA in the segmentation task of medical images to prevent forgetting. In contrast, PETAL (Brahma and Rai, 2023) utilizes the Fisher Information Matrix (FIM) as a metric of parameter importance to reset only the most irrelevant parameters. Nevertheless, these two methods are sub-optimal, as the randomness might lead to losing essential information, and pure data-driven restoration may retain false active parameters resulting from noise.
Moreover, the mean-teacher framework (Tarvainen and Valpola, 2017) serves as a base architecture for most CTTA works (Wang et al., 2022; Gong et al., 2022; Döbler et al., 2023; Niloy et al., 2024), where the teacher commonly generates pseudo-labels via a fixed threshold to supervise the student. Nonetheless, these methods suffer from low-quality pseudo-labels with a uniform threshold since the model confidence varies across categories and domains. Furthermore, Wang et al. (2024) design a dynamic thresholding technique to update the threshold based on the maximum predicted score in a batch for classification tasks, which requires a relatively large batch size to set an appropriate threshold. However, it is not suitable for object detection, where a smaller batch size is often preferred for computational efficiency. Alternatively, Gan et al. (2023) presents a cloud-device collaborative continual adaptation paradigm to accommodate dynamic environment in object detection. Nevertheless, its dependency on source data also constrains its practical utility. Therefore, a gap still exists in exploring the CTTA setting in object detection to improve the quality of pseudo-labels (Challenge 1) and to effectively reset noisy neurons (Challenge 2), without reliance on the source data.
2.4. Continual Learning
@@ 3.1.1. Problem Statement
Table 1. Comparisons between different problem settings. ‘Online’ indicates whether the model has access to the target data or only predicts the incoming test samples immediately.
Setting Source Data Target Training Data Target Distribution Train Loss Test Loss Online
Continual Learning $\times$ $(x^{t},y^{t})$ Dynamic $\mathcal{L}(x^{t},y^{t})$$\times$ $\times$
Unsupervised Domain Adaptation $(x^{s},y^{s})$ $(x^{t})$ Static $\mathcal{L}(x^{s},y^{s})+\mathcal{L}(x^{s},x^{t})$$\times$ $\times$
Source-free Domain Adaptation $\times$ $(x^{t})$ Static $\mathcal{L}(x^{t})$$\times$ $\times$
Test-time Adaptation $\times$ $\times$ Static $\times$ $\mathcal{L}(x^{t})$checkmark
Continual Test-Time Adaptation $\times$ $\times$ Dynamic $\times$ $\mathcal{L}(x^{t})$checkmark
Given a sequence of domain $D=\{D^{0},D^{1},...,D^{n}\}$, we define the first domain $D^{0}$ as the source domain and the subsequent domains as the target domain. The objective of CTTA is to enhance the performance of the model $f_{\theta_{0}}(x)$, where the parameters $\theta_{0}$ are pre-trained on source data $x^{D^{0}},y^{D^{0}}$ from $D^{0}$, in a continually changing target domain during inference time without using source data. For simplicity, we denote the data without the superscript about the specific domain in the following discussion. At time step $t$, unlabeled target data $x_{t}$ is provided sequentially, following the domain group order. The model is required to make a prediction $\hat{y}_{t}=f_{\theta_{t-1}}(x_{t})$ using the parameters $\theta_{t-1}$ which have been updated based on previous target data $x_{1},...,x_{t-1}$. Subsequently, $\hat{y}_{t}$ serves as the evaluation output at time step t, and the model will adapt itself toward $x_{t}$ as $\theta_{t-1}\rightarrow\theta_{t}$, which will only influence future inputs $x_{t+n}$.
Moreover, We compare CTTA with other adaptation settings, as detailed in Table 1. These settings are developed to meet the diverse prerequisites and requirements for real-world applications.

### v3

@@ Abstract.
Real-world application models are commonly deployed in dynamic environments, where the target domain distribution undergoes temporal changes. Continual Test-Time Adaptation (CTTA) has recently emerged as a promising technique to gradually adapt a source-trained model to continually changing target domains. Despite recent advancements in addressing CTTA, two critical issues remain: 1) Fixed thresholds for pseudo-labeling in existing methodologies generate low-quality pseudo-labels, as model confidence varies across categories and domains; 2) Stochastic parameter restoration methods for mitigating catastrophic forgetting fail to effectively preserve critical information due to their intrinsic randomness. To tackle these challenges for detection models in CTTA scenarios, we present CTAOD, featuring three core components. Firstly, the object-level contrastive learning module extracts object-level features for contrastive learning to refine the feature representation in the target domain. Secondly, the adaptive monitoring module dynamically skips unnecessary adaptation and updates the category-specific threshold based on predicted confidence scores to enable efficiency and improve the quality of pseudo-labels. Lastly, the data-driven stochastic restoration mechanism selectively reset inactive parameters with higher possibilities, ensuring the retention of essential knowledge. We demonstrate the effectiveness of CTAOD on four CTTA object detection tasks, where CTAOD outperforms existing methods, especially achieving a 3.2 mAP improvement and a 20% increase in efficiency on the Cityscapes-to-Cityscapes-C CTTA task. The code will be released.
@@ 1. Introduction
Figure 1. Motivation. The source-trained model adapts to dynamic environments without source data. (1) AM: The bar chart shows mean model scores varying across categories and domains. Unlike existing methods that continually adapt to all test data using fixed thresholds, CTAOD dynamically skips unnecessary adaptations and updates the category-specific threshold. (2) DSR: The bottom illustrates the neuron reset process, where darker purple indicates higher activity, and blue represents source knowledge. While existing methods restore parameters randomly, CTAOD resets inactive parameters with higher possibilities.
@@
Deep learning models have demonstrated immense potential across various data modalities such as image (He et al., 2016; Dosovitskiy et al., 2020), audio (Aytar et al., 2016), and language (Devlin et al., 2018). However, these models experience pronounced degradation in performance when confronted with training data (i.e., source domain) and testing data (i.e., target domain) originating from disparate distributions. This phenomenon, commonly referred to as distribution shifts, poses a significant challenge (Geirhos et al., 2018; Hendrycks and Dietterich, 2019).
In such a scenario, unsupervised domain adaptation (UDA) becomes crucial, which typically involves aligning the source and target data distributions, thereby mitigating the impact of distribution shifts (Ganin et al., 2016; Tzeng et al., 2017). Still, UDA falls short by necessitating access to source data, which is often inaccessible due to privacy constraints or data transmission barriers (Huang et al., 2021).
@@
This limitation catalyzes the exploration of Test-Time Adaptation (TTA) where the source-trained model directly adapts toward unlabeled test samples encountered during evaluation in an online manner, without the reliance on the source data (Wang et al., 2020). Nonetheless, the TTA methods, which assume a static target domain, face a more challenging and realistic problem, as real-world systems work in non-stationary environments. For example, a vehicle may encounter various continuous environmental changes such as fog, night, rain, and snow during its journey. Existing TTA methods are vulnerable to catastrophic forgetting of source knowledge and error accumulation when adaptation faces more than one distribution shift (Wang et al., 2022).
@@
Recently, Wang et al. (2022) introduce CoTTA by applying stochastic parameters restoration to mitigate catastrophic forgetting in Continual Test-Time Adaptation (CTTA) scenarios, where the model is continually adapted to sequences of target domains. Although randomly resetting parameters partially helps alleviate the forgetting of source knowledge (Wang et al., 2022; Zhu et al., 2023), its randomness may also contribute to losing crucial knowledge specific to the current domain (refer to Figure 1). Additionally, existing CTTA methods normally utilize pseudo-labeling through a fixed threshold for self-supervision (Wang et al., 2022; Döbler et al., 2023). Given that model confidence may vary across categories and domains (refer to Figure 1), employing a uniformly fixed threshold could exclude high-quality pseudo-labels while incorporating incorrect ones. Furthermore, these methods adapt to every new incoming data, resulting in computational inefficiency and potential performance degradation. Low-quality pseudo-labels and unstable adaptation to noisy data lead to error accumulation, as they provide negative feedback to the model.
@@
Therefore, despite recent efforts to handle CTTA, two significant challenges persist: Current self-training-based methods suffer from low-quality pseudo-labels, leading to error accumulation (Challenge 1); Continual adaptation to dynamic environments struggles to effectively retain valuable knowledge about the current domain while mitigating the forgetting of source knowledge (Challenge 2).
@@
To tackle these challenges, we introduce CTAOD (Continual Test-time Adaption for Object Detection). Aligning with previous CTTA works for robust adaptation (Wang et al., 2022), CTAOD is constructed upon the mean-teacher framework (Tarvainen and Valpola, 2017) where the teacher model is an exponential moving average of the student model. Particularly, CTAOD comprises Object-level Contrastive Learning (OCL), Adaptive Monitoring (AM), and Data-driven Stochastic Restoration (DSR).
@@
More specifically, OCL extracts features based on proposals generated by the Region Proposal Network (RPN), which provides multiple cropped views around the object at different locations and scales. Subsequently, Contrastive Learning (CL) loss is applied on RPN cropped views to guide the model to encourage similar object instances to remain close while pushing dissimilar ones apart. The OCL is well integrated into the mean-teacher paradigm as a drop-in enhancement to acquire more fine-grained and localized feature representation. Furthermore, to make adaptation more efficient and stable and to improve the quality of pseudo-labels, i.e., addressing Challenge 1, we design an AM module to decide whether to pause or resume the adaptation and dynamically adjust the category-specific thresholds, based on the predicted confidence scores. The dynamic nature of the AM method makes it better suited to address the effects of continuously changing distributions. Finally, the DSR mechanism is proposed to reset inactive parameters with a higher possibility than active ones by employing the gradients of parameters as weights for a random mask matrix. DSR not only helps prevent forgetting but preserves important information, i.e., addressing Challenge 2. On the other hand, its randomness also allows potentially falsely activated parameters to be reset, thereby making the model more stable.
@@
We demonstrate the effectiveness of our proposed approach on four continual test-time adaptation benchmark tasks for object detection. These datasets are associated with synthetic and real-world distribution shifts in the short-term and long-term adaption, i.e. Cityscapes (Cordts et al., 2016), Cityscapes-C (Hendrycks and Dietterich, 2019), SHIFT (Sun et al., 2022), and ACDC (Sakaridis et al., 2021). Experimental results indicate that our method significantly improves performance over existing methods, with gains of up to 3.2 mAP and 20% in computational efficiency.
Our main contributions are summarized as follows:
@@
  • This study introduces CTAOD, which pioneers exploring CTTA for detection models. Specifically, we propose to leverage object-level features for contrastive learning to refine feature representation in CTTA object detection.
@@
  • To address the two challenges in CTTA, our proposed CTAOD enables dynamic skipping and category-specific threshold updates based on predicted scores and reset the inactive parameter with higher possibilities.
@@
  • Empirical experiments demonstrate that CTAOD surpasses existing methods and facilitates efficient short-term and long-term adaptation in dynamic environments.
@@ 2.1. Source-Free Domain Adaptation
UDA tackles the inter-domain divergence by aligning the distributions of source and target data (Ganin et al., 2016; Luo et al., 2020). Despite its effectiveness, the limitation of UDA lies in its requirement for access to the source domain data, which often raises concerns regarding data privacy and transmission efficiency. As a result, Source-Free Domain Adaptation (SFDA) received extensive research attention in object detection, where the source-trained detector is re-trained on the target training data (VS et al., 2023b, a; Li et al., 2021) before evaluation. For instance, MemCLR (VS et al., 2023a) employs a cross-attention-based memory bank for CL, while IRG (VS et al., 2023a) utilizes the object relations with instance relation graph network to explore the SFDA setting for object detection without the source data. However, the standard SFDA setting requires prior knowledge of the target domain, which is impractical in most real-world applications.
@@ 2.2. Test-Time Adaptation
TTA adapts the source-trained model to the target test data during inference time without access to the source data. Since both TTA and SFDA involve adapting the source-trained model to the unlabeled target data without utilizing source data, some works also refer to TTA as SFDA (Wang et al., 2022; Brahma and Rai, 2023). In this paper, we distinguish between TTA and SFDA based on evaluation protocol, although they can be transformed into each other in experiments. Furthermore, TTA methods improve the model performance under distribution shift commonly through pseudo-labeling (Sinha et al., 2023; Iwasawa and Matsuo, 2021), batchnorm statistics updating (Hu et al., 2021; You et al., 2021), or entropy regularization (Wang et al., 2020; Iwasawa and Matsuo, 2021) during testing. For example, Tent (Wang et al., 2020) updates the batchnorm parameters with entropy minimization and demands a large batch size for optimization during test-time adaptation, which is unsuitable for real-time detection model deployment where images are processed sequentially. TTA methods assume a static target domain where the target data come from a single domain. However, the target distribution may exhibit a continual shift over time in practical scenarios.
@@ 2.3. Continual Test-time Adaptation
Recently, Wang et al. (2022) introduce CoTTA, marking the first work tailored to the demands of CTTA by adapting a pre-trained model to sequences of domains without source data. Subsequently, research efforts have been dedicated to exploring CTTA, primarily focused on classification (Döbler et al., 2023; Brahma and Rai, 2023; gan2023decorate) and segmentation (Song et al., 2023; Ni et al., 2023; Liu et al., 2023) tasks. For instance, similar to CoTTA, Zhu et al. (2023) apply a stochastic reset mechanism to prevent forgetting in the CTTA medical segmentation task. In contrast, PETAL (Brahma and Rai, 2023) utilizes the Fisher Information Matrix (FIM) as a metric of parameter importance to reset only the most irrelevant parameters. Nevertheless, these two methods are sub-optimal, as the randomness might lead to losing essential information, and pure data-driven restoration may retain false active parameters resulting from noise.
Moreover, the mean-teacher framework (Tarvainen and Valpola, 2017) serves as a base architecture for most CTTA works (Wang et al., 2022; Gong et al., 2022), where the teacher generates pseudo-labels via a fixed threshold to supervise the student. Nonetheless, these methods suffer from low-quality pseudo-labels with a uniform threshold since the model confidence varies across categories and domains. Furthermore, Wang et al. (2024) design a dynamic thresholding technique to update the threshold in a batch for classification tasks, requiring a large batch size. However, this approach is unsuitable for object detection where smaller batch sizes are preferred for computational efficiency. To improve efficiency, yoo2024and presents a CTTA approach for object detection to update specific parts of the model in specific adaptation, which utilizes source statistics. Although it effectively improved efficiency, its performance is limited. Additionally, Gan et al. (2023) presents a cloud-device collaborative continual adaptation paradigm in object detection by aligning source and target distribution. Nevertheless, the reliance on source data in both methods restricts their practical applicability. Therefore, a gap still exists in exploring the CTTA in object detection to improve the quality of pseudo-label (Challenge 1) and effectively reset noisy neurons (Challenge 2), without relying on any source data.
@@ 3.1.1. Problem Statement
Given a sequence of domain $D=\{D^{0},D^{1},...,D^{n}\}$, we define $D^{0}$ as the source domain and the subsequent domains as the target domain. The objective of CTTA is to enhance the performance of the model $f_{\theta_{0}}(x)$, where the parameters $\theta_{0}$ are pre-trained on source data $x^{D^{0}},y^{D^{0}}$ from $D^{0}$, in a continually changing target domain during inference time without using source data. For simplicity, we denote the data without the superscript about the specific domain in the following discussion. At time step $t$, unlabeled target data $x_{t}$ is provided sequentially, following the domain group order. The model is required to make a prediction $\hat{y}_{t}=f_{\theta_{t-1}}(x_{t})$ using the parameters $\theta_{t-1}$ which have been updated based on previous target data $x_{1},...,x_{t-1}$. Subsequently, $\hat{y}_{t}$ serves as the evaluation output at time step t, and the model will adapt itself toward $x_{t}$ as $\theta_{t-1}\rightarrow\theta_{t}$, which will only influence future inputs $x_{t+n}$.
Figure 2. Overview. CTAOD follows the mean-teacher framework, featuring: (1) The OCL module compares region of interest features extracted from feature maps of both networks based on teacher proposals for contrastive learning; (2) The AM module dynamically skips unnecessary adaptation and adjusts category-specific thresholds based on prediction scores; (3) The DSR reset the inactive parameter with higher possibilities based on the FIM.